What is the solution to x+4/9+x+2/2=2
Answer by rfer(16322)

2x+4/9+1=2
2x+4/9+9/9=2
2x+13/9=2
2x=2-13/9
2x=18/9-13/9
2x=5/9
x=(5/9)/(2/1)
x=5/9*1/2
x=5/18